State and prove Least Upper Bound (LUB) property of Real Numbers.
1
Expert's answer
2021-10-04T16:20:37-0400

Statement;

Let S be a non empty set of real numbers.

A real number x is the least upper bound for S if x is an upper bound for S and x "\le" y for every upper bound y of S.

Proof;

The least-upper-bound property can be proven using the assumption that every Cauchy sequence of real numbers converges. Let S be a nonempty set of real numbers. So if S has exactly one element, then its only element is a least upper bound. So consider S with more than one element, and suppose that S has an upper bound B1. Since S is nonempty and has more than one element, there exists a real number A1 that is not an upper bound for S. Define sequences A1, A2, A3, ... and B1, B2, B3, ... recursively as follows:

(i) Check whether "\frac{(A_n+B_n)}{2}" is an upper bound for S

(ii) If it is ,let An+1=An and let Bn+1="\frac{(A_n+B_n)}{2}"

(iii) Otherwise there must be an element s in S such that "s>\frac{(A_n+B_n)}{2}."

Let An+1=s and Bn+1= Bn.

Then A1 "\le" A2 "\le" A3 "\le""\le" B3 "\le" B2 "\le" B1 and |An − Bn| → 0 as n → ∞.

It follows that both sequences are Cauchy and have the same limit L, which must be the least upper bound for S.